Light of wavelength 180nm ejects photoelectrons from a plate of metal whose work function is 2eV.If a uniform magnetic field of 5/100000 tesla be applied parallel to the plate, the radius of path followed by elesctrons ejected normally from the plate with maximum energy would be

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

highest energy of photoelectrons=(hc/180nm)-3.2*10-19=E, v=√(2E/m), mv2/r=evB, solve for 'v'